If P(E1)=P(E2)=0.15, P(E3)=0.4, and P(E4)=2P(E5), calculate the P (E4) and P(E5).

1
Expert's answer
2022-05-17T23:32:24-0400

We know that total probability =1

i.e. sum of probabilities "P(E_1)+P(E_2)+P(E_3)+P(E_4)+P(E_5)=1"

So, "0.15+0.15+0.4+2x+x=1"

where "x=P(E_5)"

Solving for "x" we get:

"3x=1-0.7"

"3x=0.3"

"x=0.1"

So, "P(E_5)=0.1" and "P(E_4)=2*0.1=0.2"